# Dzyalonshinskii-Moriya interaction in Fe5GeTe2 epitaxial thin films

### v1

@@ 4 Conclusion
In conclusion, we have used BLS to quantitatively measure a significant DMI in several epitaxial Fe5GeTe2 films with nominally symmetric interfaces. The DMI values were consistent across films of different thicknesses and exhibited the same sign, indicating a bulk origin for the DMI. Although our measurements cannot reveal the microscopic origin of the DMI, they are consistent with a partial ordering of the Fe split sites during film growth, a mechanism supported by published ab initio calculations showing that such ordering induces DMI.
These results confirm that Fe5GeTe2 can host a significant DMI and suggest that the reported magnetic textures in Fe5GeTe2 are likely chiral in nature. If the mechanism for the non-centrosymmetric structure is indeed induced during the growth as we suggest, the DMI may be enhanced by adjusting the growth conditions. Additionally, DMI might be stronger in Fe5GeTe2 films synthesized or processed using other techniques.
The bulk DMI mechanism observed here may coexist with interfacial contributions, such as those involving oxide layers as considered in previous studies, offering opportunities to further enhance the DMI.
@@
5 Methods
5.1 Determination of magnetic parameters from BLS spectroscopy
BLS spectroscopy was used to analyse the spinwave frequency as a function of field or wavevector. A green laser beam ($\lambda=532$ nm) was directed onto the sample at an angle $\phi$ (Fig. 3b), with the sample placed inside a cryostat at a controlled temperature. A magnetic field was applied using permanent magnets to align the equilibrium magnetization in the film plane and perpendicular to the incident light, following the Damon-Eshbach geometry. The backscattered light was analysed with a double, three-pass Fabry-Perot spectrometer (model TFP2-HC from JRS Instruments). Some of the light is backscattered by spinwaves, with a wave vector defined by the incidence angle as $k=4\pi\sin\phi/\lambda$. This produces two peaks in the spectrum: a redshifted (Stokes) peak corresponding to forward-propagating spin waves, and a blueshifted (anti-Stokes) peak corresponding to back-propagating spin waves. The incident and analysed light are cross-polarised, which suppresses most peaks arising from phonon scattering. The observed peaks were characterized by their average frequency ($f_{0}$), their frequency difference ($\Delta f$), and their full width at half maximum ($\Gamma$), all obtained by fitting the peaks with a Lorentzian profile. The variation of these frequencies as a function of field or wavevector can be used to determine static ($D$, $B_{K}$, $M_{s}$, …) and dynamic parameters ($g$ and $\alpha$).
The effective anisotropy field $B_{K}$ and g-factor $g$ are extracted from the variation with field of the mean frequency of the Stokes and anti-Stokes peak, $f_{0}$, by using the law of dispersion of magnetostatic spinwaves determined by Kalinikos & Slavin 28, 14:
$f_{0}=\frac{\gamma}{2\pi}\sqrt{B+\mu_{0}M_{s}P+\frac{2A}{M_{s}}k^{2}}\sqrt{(B-B_{K})-\mu_{0}M_{s}P+\frac{2A}{M_{s}}k^{2}},$ (2)
where $P\approx kt/2$ for thin films ($kt\ll 1$) and $A$ is the exchange stiffness (which is neglected in our analysis).
Figure 4: a) BLS spectra of the 8 nm film, at different applied fields (150 K, $\phi=30$º). The central peak corresponds to unscattered light and is used to evaluate the instrumental linewidth. b) Spin-wave frequency versus applied field for the three film thicknesses (150 K) ($\phi=20$º for the 3 and 12 nm films, 30º for the 8 nm). The lines are fits using the model of Kalinikos and Slavin 28, neglecting the contribution of the exchange interaction.
Fig. 4a shows BLS spectra of the 8-nm sample taken at different applied fields $B$. All spectra clearly show the Stokes and anti-Stokes peaks, which are of different amplitude, typical of magnon BLS (but not of phonons). The mean frequency $f_{0}$ is shown versus $B$ in Fig. 4b, and is fitted using eq. 2 (lines in the plot.) The determined $B_{K}$ and $g$ parameters using this method are shown in Tab. 1 for all three films at 150 K.
5.2 Determination of the Gilbert dissipation parameter
Some information on the Gilbert dissipation parameter $\alpha$ can be obtained from the peak width $\Gamma$. The main contributions to $\Gamma$ are the instrumental resolution ($\Gamma_{0}$), the magnetic dissipation, and the inhomogeneity of the magnetic properties such as the anisotropy ($\delta B$):
$\Gamma=\Gamma_{0}+\gamma\alpha(2B-B_{K})+\gamma\delta B\ F(B,B_{K}),$ (3)
where $F()$ is a function that tends towards a constant for $B\gg B_{K}$ (see SI of ref. 29). The measured $\Gamma$ vs. $B$ is shown in Fig. 5 for the three films.
Figure 5: Peak FWHM $\Gamma$ versus field $B$ for the three films. The grey lines are the expected variation assuming different values of $\alpha_{\rm eff}$ (see text for details on the calculation of $\alpha_{\rm eff}$).
The instrumental width $\Gamma_{0}$ is measured from the elastic peak at $f=0$ (seen in Fig. 4a). The inhomogeneity contribution $\delta B$ may be disentangled from the dissipation by applying a large field such that $\alpha(2B-B_{K})\gg\delta B\ F(B,B_{K})$, however this was not possible with our setup. An upper bound of the dissipation parameter, $\alpha_{\rm eff}$, can be determined by neglecting the contribution from inhomogeneity:
$\alpha_{\rm eff}=(\Gamma-\Gamma_{0})/(\gamma(2B-B_{K}))\geq\alpha$ (4)
The calculated $\alpha_{\rm eff}$ is shown as grey lines in fig. 5.
@@ 5.2 Determination of the Gilbert dissipation parameter
The authors thank V. Jeudy for his support with cryogenic measurements. Magnetometry measurements were kindly performed by E. Rivière at ICMMO (CNRS) and also at the Physical Measurements Platform at the LPS. This work was supported by the French National Research Agency (ANR) through the projects ELMAX (ANR-20-CE24-0015), NEXT (ANR-23-CE09-0034), the FLAG-ERA grant MNEMOSYN (ANR-21-GRF1-0005-01), and the LANEF framework for mutualized infrastructure (ANR-10-LABX-51-01).

### v2

@@ 3 Conclusion
In conclusion, we have used BLS to quantitatively measure a significant DMI favouring cycloidal chiral textures in several epitaxial Fe5GeTe2 films with nominally symmetric interfaces. The DMI values were consistent across films of different thicknesses and exhibited the same sign, indicating a bulk origin for the DMI. Although our measurements cannot reveal the microscopic origin of the DMI, they are consistent with a partial ordering of the Fe split sites during film growth, a mechanism supported by published ab initio calculations showing that such ordering induces DMI.
These results confirm that Fe5GeTe2 can host a significant DMI and suggest that the reported magnetic textures in Fe5GeTe2 are likely chiral in nature. If the mechanism for the non-centrosymmetric structure is indeed induced during the growth as we suggest, the DMI may be enhanced by adjusting the growth conditions or by employing chemical substitutions which preferentially affect the split-site sublattice, such as with Co 36 or Ni 37. Additionally, DMI might be stronger in Fe5GeTe2 films synthesized or processed using other techniques. The bulk DMI mechanism observed here may coexist with interfacial contributions, such as those involving oxide layers as considered in previous studies, offering opportunities to further enhance the DMI.
@@
{suppinfo}
The supporting information (S.I.) includes more details and figures relating to the experimental methods, the structural analysis of the films, the fitting and analysis of the BLS spectra, additional BLS measurements, and micromagnetic simulations of the effects of the measured DMI.
@@
The authors thank V. Jeudy for his support with cryogenic measurements and H. Okuno for the STEM image of Fig. S1. Magnetometry measurements were kindly performed by E. Rivière at ICMMO (CNRS) and at the Physical Measurements Platform at the LPS. This work was supported by the French National Research Agency (ANR) through the projects ELMAX (ANR-20-CE24-0015), ESR/EQUIPEX+ 2D-MAG (ANR-21-ESRE0025), NEXT (ANR-23-CE09-0034), the FLAG-ERA grant MNEMOSYN (ANR-21-GRF1-0005-01), the LANEF framework for mutualised infrastructure (ANR-10-LABX-51-01), and the France 2030 government investment plan (PEPR SPIN: SPINMAT ANR-22-EXSP-0007 and CHIREX ANR-22-EXSP-0002).